OpenStudy (anonymous):

can someone please help me with factor 27x^3-8y^3 i no 9 and 3 is 27 and -4 and 2 is -8

OpenStudy (asnaseer):

ok - I assume you know how to factorise this expression:\[a^3-b^3\]

OpenStudy (anonymous):

= (a+b) (a^2-ab+b^2)

OpenStudy (asnaseer):

not quite, it should be:\[a^3-b^3=(a-b)(a^2+ab+b^2)\]

OpenStudy (asnaseer):

so now you just need to express your equation as the difference of two cubes

OpenStudy (asnaseer):

your first term is \(27x^3\) - how can this be expressed as (something)^3 ?

OpenStudy (anonymous):

3^3 and 2^3

OpenStudy (asnaseer):

? \[27x^3=(???)^3\]what should the ??? be replaced by?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

3x

OpenStudy (asnaseer):

correct, so the first term can be expressed as \((3x)^3\) do the same for the second term - what do you get?

OpenStudy (anonymous):

(2y)^3

OpenStudy (asnaseer):

perfect, so your original expression can be written as:\[27x^3-8y^3=(3x)^3-(2y)^3\]now it is in the form where you can make use of the factorisation:\[a^3-b^3=(a-b)(a^2+ab+b^2)\]

OpenStudy (anonymous):

3^3-2y^3= (3-2) (3^2+3*4+4^2)

OpenStudy (asnaseer):

???

OpenStudy (asnaseer):

what did you use to represent 'a' and 'b' ?

OpenStudy (anonymous):

3 and 2

OpenStudy (asnaseer):

remember we ended up with:\[27x^3-8y^3=(3x)^3-(2y)^3\]and we know:\[a^3-b^3=(a)^3-(b)^3=(a-b)(a^2+ab+b^2)\]so look carefully at what 'a' and 'b' should be.

OpenStudy (anonymous):

a is 3 and 2 is b

OpenStudy (asnaseer):

what expression is inside the braces for the first term in: \((3x)^3-(2y)^3\) ?

OpenStudy (anonymous):

3x and 2y

OpenStudy (asnaseer):

well - first term is 3x. so we know 'a' must be '3x' similarly, 2nd term is '2y'. so we know 'b' must be '2y'

OpenStudy (asnaseer):

do you understand?

OpenStudy (asnaseer):